A particle is thrown upward from the ground. It experiences a constant air resistance which can produce a retardation of 2m/sec2 opposite to the direction of the velocity of a particle. The ratio of time of ascent to the time of descent  is {take g=10m/sec2}

Grade:11

1 Answers

Gowtham
6 years ago
time of flight is 2h/g
time of ascent is h/g and time of descent is h/g for a normal journey
 
if air resistant is considered
during time of ascent both resistance and gravity acts downwards
so time of ascent is h/(g+a)
durind descent the air resistance acts upwards and gravity acts downwards
so time of descent is h/(g-a)
so
the time of ascent/time of descent is g-a/g+a
so ta/td=10-2/10+2
            =8/12
            =2/3